OpenStudy (falco276):

Someone help please? :( What is the equation in point-slope form of the line passing through (−2, 1) and (4, 13)? y − 1 = −2(x − 4) y − 1 = 2(x + 2) y − 13 = −2(x + 2) y − 13 = 2(x + 4)

jigglypuff314 (jigglypuff314):

first find the slope \[slope = m = \frac{ rise }{ run } = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] when given the points (x1, y1) and (x2, y2) then point-slope form is: y - y1 = m(x - x1) given the point (x1, y1) and m = slope

OpenStudy (falco276):

But, what's the answer? :( I'm kinda confused by these formulas.

OpenStudy (falco276):

I honestly suck at Algebra 2

jigglypuff314 (jigglypuff314):

well, first find the slope, tell me what you get

jigglypuff314 (jigglypuff314):

(13 - 1) = 12 (4 - (-2)) = 6 so slope is 12/6 or 2

OpenStudy (falco276):

okay. 12/2

jigglypuff314 (jigglypuff314):

no, sorry, I meant 12/6 = 2

OpenStudy (falco276):

That's what I got using Distance formula.

jigglypuff314 (jigglypuff314):

... distance formula should not be needed in this question.

OpenStudy (falco276):

Oops, I think I meant slope formula. Sorry.

jigglypuff314 (jigglypuff314):

4 - (-2) = 6 not 2

OpenStudy (falco276):

Maybe a minor mistake by me. Sorry.

jigglypuff314 (jigglypuff314):

that's fine :) now point-slope form is: y - y1 = m(x - x1) you know a point, (−2, 1), and you know the slope is 2 so you plug in y - (1) = (2)(x - (-2))

OpenStudy (falco276):

But how do we know what x and y are? :/

jigglypuff314 (jigglypuff314):

point-slope form is: y - y1 = m(x - x1) when given the point (x1, y1) and m = slope given the point (-2 , 1 ) and m = 2

OpenStudy (falco276):

Wait, am doing it on scratch paper.

jigglypuff314 (jigglypuff314):

k :) lol I just realized that you were the one who commented on my previous profile picture :P I changed it, happy now?

OpenStudy (falco276):

yeah. :) lol. am still confused about this formula thing. -_- I mean with plugging the numbers thing. IT'S JUST...... ARRGH!! IDK!!!!!!!!!!

jigglypuff314 (jigglypuff314):

you plug in -2 for x1, 1 for y1, and 2 for m

OpenStudy (falco276):

Hmm.... -2 for x1. So does that become -2?

jigglypuff314 (jigglypuff314):

x - (-2) = x + 2

OpenStudy (falco276):

-2 - (-2)= -2+2?

OpenStudy (falco276):

0=4

OpenStudy (falco276):

Dosen't match. :/

jigglypuff314 (jigglypuff314):

no, you plug in -2 for x1 only, not x x and x1 are different

OpenStudy (falco276):

Ohhh.....okay. I think I get it now.

jigglypuff314 (jigglypuff314):

you just plug in so that y - (1) = (2)(x - (-2)) so y - 1 = 2(x + 2) and which one does that match with?

OpenStudy (falco276):

omgs..... I can't do this. :( I just don't get it. :/ Does x and y stay alone or should I replace them?

jigglypuff314 (jigglypuff314):

they stay alone, take a look at your options...

OpenStudy (falco276):

y − 1 = −2(x − 4) y − 1 = 2(x + 2) y − 13 = −2(x + 2) y − 13 = 2(x + 4)

OpenStudy (falco276):

So is it B?

jigglypuff314 (jigglypuff314):

yup :)

OpenStudy (falco276):

Stay here, don't go away. :)

OpenStudy (falco276):

I have something to say about Algebra.

OpenStudy (falco276):

To me, in my opinion, it dosen't mean that it's too hard. The thing that I don't get it is plugging in the numbers to a complicated formula. :/ And all of these numbers and letters such as X and Y and..... UGH! It's just too confusing! I find Geometry much easier.

OpenStudy (falco276):

I mean, you don't have to focus and think about numbers and letters there. There they have shapes. Yes, formulas are also included, BUT they are easy to solve. Algebra? C'mon! Whoever invented Algebra is mousebrained. >:(

jigglypuff314 (jigglypuff314):

lol you're ranting while I'm doing my math homework which is finding the slope of a fluttering curve 0.o you have a lot more pellet to deal with kid if you don't start liking algebra now.

jigglypuff314 (jigglypuff314):

OMG LOL I can't believe this site has a sensor to replace curse words

OpenStudy (falco276):

(sigh) yeah, you have to face what coming at you, you simply can't leave that behind. Fine.

OpenStudy (falco276):

REALLY!!!!!!!!

OpenStudy (falco276):

okay Imma do one word. pellet.

OpenStudy (falco276):

flutter.

jigglypuff314 (jigglypuff314):

oh yeah, I still don't get how "occultic" is bad na I tried spelling sh!t and it turned out pellet

OpenStudy (falco276):

Maybe this is supposed to be a "safe" site for everyone?

jigglypuff314 (jigglypuff314):

yeah, cuz there's apparently 4 yr olds on here 0.o now could you explain why something "occultic" is discouraged?

OpenStudy (falco276):

Okay, um how do I describe this? Okay it's everywhere. The eye, the pyramid/triangle, star, moon. It's basically used in everyday things, but the problem is that we don't realize it's bad. We just think that it's just something used in everything. BUT it's not! We have to realize that those symbols have to do with the occult. How? btw, did you see that vid I sent to you?

jigglypuff314 (jigglypuff314):

no, I have not... what's "the occult"?

OpenStudy (falco276):

http://en.wikipedia.org/wiki/Occult Check this link out and you'll get to know about it. :)

jigglypuff314 (jigglypuff314):

-.-' I'm guessing you've never read Harry Potter the whole point of the Harry Potter series is witchcraft and wizardry and magic...

OpenStudy (falco276):

Exactly! The more we use these symbols in everyday things, the more were turning towards the enemy that has the grip to this world. So IT'S NOT TOO LATE! TURN TO CHRIST but not to the Christan religion. There is a difference. The churches today are making money which is a sign by the Devil. So start reading the Bible and praying and you'll get to know the real truth. And this way, you'll spread the good news to your family and friends. :)

OpenStudy (falco276):

Here are a few verses that I want to share with you and please pass this to your family and friends :) Revelation ch. 21 verse 8

OpenStudy (falco276):

Oh also not to mention videos: http://www.youtube.com/watch?v=h0cbtkGHDIY

OpenStudy (falco276):

Please watch it, It's a testimony about how he got saved by God from practicing occultic things. There is a lot in this connection but very few are making the use of it.

jigglypuff314 (jigglypuff314):

sorry, but I don't have the time. I have no interest in participating in religions, if I burn then so be it. I can accept that yes there is a higher being, which one out of all the religions in the world? at least I'm open to options. people believe what they want. I congratulate you on such a devotion to yours :)

OpenStudy (falco276):

What I write is only for information, not an imposition. It's not enough to belive that there is a high power. If that high power is not personal then He cannot be the creator. In every person as a creation there has to be a search for this high power. For the unification of the two. You mentioned a bunch of religions which are all working against this high power or creator. They divide people for thier financial purposes and domination. Please read in the Bible the book of Romans 16: verse 17 and 18. You can see how people are cheated in the name of God. I'm not for religions. I don't preach religions. I wish you had a longing. for the creator. The loving one. :)

OpenStudy (falco276):

@jigglypuff314

jigglypuff314 (jigglypuff314):

I do have a longing, a longing to find that higher being that sends the future to my dreams. that is all. whether he be saint or sin.

OpenStudy (falco276):

Okay, farewell, it's your thing. Thanks for helping me in Algebra. :)